jennylove:

If you have a 45.0 V battery connected to a 10.0 Ωresistor and a 500.0 Ω resistor in series, what will be the total power in the circuit? 45.0 W 13.5 W 4.50 W 3.97 W

jennylove:

im not quite sure. @TheSmartOne

jennylove:

please explain, i wanna make sure i understand it! (:

TheSmartOne:

Of course! So in circuit where the resistors are connected in series, what would be the total resistance? Do we add the sum of the resistors (R1+R2) or do we have to do the sum the inverse of the resistance (1/R1 + 1/R2)

jennylove:

you would add them! (:

TheSmartOne:

Bingo! In parallel circuits we would do the sum of the inverse So what is the sum of the resistors in this series circuit?

TheSmartOne:

 What formula would we use? We want power and we have the resistance and voltage

jennylove:

we would use I=V/R ?

TheSmartOne:

That would give us current. We want the power!

jennylove:

oh im sorry i mean the P= V* I

TheSmartOne:

Once again Power = voltage * current We could calculate current from the previous formula you mentioned and then plug it in here But if you look at the chart, there is a direct formula that basically plugs in the two formulas you mentioned P = IV I = V/R Plug in the second into the first P = V^2/R You can find this one in the top left hand corner of that image I shared

TheSmartOne:

That's how you get that formula from the formula you mentioned ^

TheSmartOne:

But does that make sense? That way we don't have to calculate the current so it saves us some time and steps :)

jennylove:

ohh okay, so you combine them together! yes that does make sense. could i solve it and you can double check it?

TheSmartOne:

Absolutely! Go for it :)

jennylove:

it would be 45.0 ?

TheSmartOne:

Can you show me your work so I can see where you might have made a mistake

jennylove:

well i used 10 as the resistance but i think i may have done that wrong

TheSmartOne:

You forgot to add the 500 ohms resistance

jennylove:

you add them together right so would it be 510 ?

TheSmartOne:

Yup!

jennylove:



jennylove:

yes! i think i got it

TheSmartOne:

Well done!!

jennylove:

thank you so much! seriously you have been very helpful. I have a few more questions, i you may check them for me?

jennylove:

if you*

TheSmartOne:

For sure, anytime!